A 40.7g of a non-electrolyte is dissolved in 220g of water to make a solution. The solution's freezing point is -5.53°C. With the given data, what is the molar mass of the solute?


1
Expert's answer
2021-04-28T02:46:08-0400

Molar Mass of H2O = 18.0153

= 220/18.0153 = 12.212moles

Molality = 12.212/-5.53

= -2.21 × 40.7

= 89.88g/mol